Market Ratios Dec 31, 2018 Price to earnings (P/E) Price to operating profit (P/OP) 30.47 22.54 6.15 Price to sales (P/S) PriCompare and analyze ratios of two companies.
Answer #1

P/E Ratio = Market Price per share/Earnings per share

It indicates the number of times the share is priced as compare to its earnings. It also means the number of years (provided, earnings are constant) it will take to return the investment i.e. payback period.

P/OP Ratio = Market Price/Operating Profit i.e. EBIT per share

It is same as P/E Ratio, except that, instead of EPS, EBIT per share is used.

P/S Ratio = Market Price/Sales per share

It indicates that how much times of market price is sales. It is useful in valuing cyclical companies where profits are cyclical and also companies that are in early growth stages and not yet profitable.

P/BV Ratio = Market Price per share/Book Value per share

It indicates the number of times the market price is compared to its book value i.e. net assets per share. It is useful to compare companies with similar growth and profitability.

Analysis:

Assumption: Both companies are of same industry

(1) PE Ratio of B is lower than A which simply means that B is undervalued compared to A and it is a better investment opportunity, taking only earnings into consideration.

(2) P/OP Ratio of B is again lower than A which means that B is undervalued. But comparing it to PE Ratio also indicates that, (a) P/OP ratio of A is lower than PE ratio which means that A has significant amount of other incomes which leads to higher earnings, also it may be having very less debt and lower tax expense (b) P/OP ratio is higher than PE Ratio which means that it may have higher debt and tax expense might also be higher.

(3) There is a huge difference between P/S ratio of both companies which makes B a very attractive investment opportunity. First of all, it indicates that B is highly undervalued, second it might be in its early stage and has lower sales which means there are chances of high growth.

(4) P/BV ratio of both companies are almost same which makes neither of them better than the other.

Summary: Lower PE Ratio, P/OP Ratio & P/S Ratio indicates that B is highly undervalued and it is a better investment opportunity compared to A.